OpenStudy (anonymous):

The line that contains the point Q( 1, -2) and is parallel to the line whose equation is y - 4 = 2/3 (x - 3) Indicate the equation of the given line in standard form. The line that is the perpendicular bisector of the segment whose endpoints are R(-1, 6) and S(5, 5) help

OpenStudy (shamim):

so u hv 2 questions. is not it

OpenStudy (shamim):

for ur 1st math

OpenStudy (shamim):

ur given equation of line is\[y-4=\frac{ 2 }{ 3 }(x-3)\]

OpenStudy (shamim):

\[y=\frac{ 2 }{ 3 }x-2+4\]

OpenStudy (shamim):

\[y=\frac{ 2 }{ 3 }x+2\]

OpenStudy (shamim):

now u hv to compare this given equation of line with standard equation of line which is \[y=mx+c\]

OpenStudy (shamim):

so the slope of ur given line is\[m=\frac{ 2 }{ 3 }\]

OpenStudy (shamim):

slope of any parralel line is same

OpenStudy (shamim):

so slope of a parallel line will b\[m=\frac{ 2 }{ 3 }\]

OpenStudy (shamim):

this parallel line will go through a point (1,-2)

OpenStudy (shamim):

u know the equation of line is\[y-y _{1}=m(x-x _{1})\]

OpenStudy (shamim):

here\[x _{1}=1,y _{1}=-2\]

OpenStudy (shamim):

so the equation of ur parallel line is\[y-(-2)=\frac{ 2 }{ 3 }(x-1)\]

OpenStudy (shamim):

now try to solve it

OpenStudy (shamim):

did u get it

OpenStudy (anonymous):

im solving it :)

OpenStudy (shamim):

ok

OpenStudy (shamim):

write the result

OpenStudy (anonymous):

well im stuck y-2=2/3x-2/3

OpenStudy (shamim):

no

OpenStudy (shamim):

\[y+2=\frac{ 2 }{ 3 }(x-1)\]

OpenStudy (shamim):

\[3y+6=2x-2\]

OpenStudy (shamim):

do the rest

OpenStudy (anonymous):

3y=2x-8 or 2x-3y=8

OpenStudy (shamim):

if u feel any difficulties then ask me please

OpenStudy (shamim):

ya u r correct

OpenStudy (anonymous):

i am :D yay :)

OpenStudy (anonymous):

2x-3y=8 so its that

OpenStudy (shamim):

ya

OpenStudy (anonymous):

thank you

OpenStudy (shamim):

welcome

OpenStudy (shamim):

bt u hv another math here

OpenStudy (anonymous):

oh yeah

OpenStudy (shamim):

2nd portion of ur math

OpenStudy (shamim):